Skip to content

Commit 8c5810c

Browse files
authored
Merge branch 'master' into user_guide_updates_contacts
2 parents 8a7530e + c3b21c9 commit 8c5810c

File tree

2 files changed

+20
-2
lines changed

2 files changed

+20
-2
lines changed

docs/logbook/LHC/2025_lhc.md

Lines changed: 19 additions & 2 deletions
Original file line numberDiff line numberDiff line change
@@ -31,6 +31,7 @@ The tables below can be sorted by clicking next to the column headers.
3131
| 2025-04-24 13:00 | 2025-04-25 03:30 | 0.6W 1WN | Commissioning | Global Corrs at 60cm and 18cm and Validation | [Shift Start](logbook://2025-04-24,1081,4243122) / [Summary](logbook://2025-04-25,1081,4243700) ([OP][op_global_corrs_60cm_18cm]{.cern_login target=_blank}) |
3232
| 2025-04-27 07:30 | 2025-04-27 23:30 | 1H 1HN | Commissioning | Validation of Everything | [Shift Start](logbook://2025-04-27,1081,4245465) / [Summary](logbook://2025-04-27,1081,4245531) ([OP][op_validation_everything]{.cern_login target=_blank}) |
3333
| 2025-05-07 20:00 | 2025-05-08 07:30 | 1.5WN | Commissioning | Pb Ion Optics Part 1 | [Shift Start](logbook://2025-05-07,1081,4257152) / [Summary](logbook://2025-05-08,1081,4257277) ([OP][op_ion_optics_part1]{.cern_login target=_blank}) |
34+
| 2025-05-14 07:30 | 2025-05-14 15:00 | 1.0W | Commissioning | VdM optics commissioning | [Shift Start](logbook://2025-05-14,1081,4292027) / [Summary](logbook://2025-05-14,1081,4292225) ([OP][op_vdm]{.cern_login target=_blank}) |
3435
| 2025-05-18 13:00 | 2025-05-18 19:00 | 0.5H 0.25HN | Commissioning | Pb Ion Optics Part 2 | [Shift Start](logbook://2025-05-18,1081,4295338) / [Summary](logbook://2025-05-18,1081,4295555) ([OP][op_ion_optics_part2]{.cern_login target=_blank}) |
3536
| 2025-06-12 13:30 | 2025-06-12 22:30 | 0.5W 0.6WN | Commissioning | Pb/pO Ion Optics | [Shift Start](logbook://2025-06-12,1081,4314767) / [Summary](logbook://2025-06-12,1081,4383876) ([OP][op_po]{.cern_login target=_blank}) |
3637
| 2025-06-13 16:30 | 2025-06-13 20:00 | 0.2W 0.3WN | Commissioning | Hierarchy breaking investigation at Injection | [Shift Start](logbook://2025-06-13,1081,4315743) / [Summary](logbook://2025-06-13,1081,4315800) ([OP][op_hierarchy_breaking_injection]{.cern_login target=_blank}) |
@@ -43,7 +44,14 @@ The tables below can be sorted by clicking next to the column headers.
4344
| 2025-09-03 00:00 | 2025-09-03 04:00 | 0.5WN | MD15003 | Valiation of combined ramp/squeeze/rotate | [Shift Start](logbook://2025-09-03,1081,4377865) / [Summary](logbook://2025-09-03,1081,4378088) ([OP][op_MD15003]{.cern_login target=_blank}) |
4445
| 2025-09-04 23:00 | 2025-09-05 07:45 | 1.1WN | MD15464 | Refining optics corrections at top energy | [Shift Start](logbook://2025-09-04,1081,4380286) / [Summary](logbook://2025-09-05,1081,4384687) ([OP][op_MD15464]{.cern_login target=_blank}) |
4546
| 2025-09-08 04:00 | 2025-09-08 08:30 | 0.2HN | Commissioning | Pb Ion Optics Part 3 | [Shift Start](logbook://2025-09-08,1081,4382873) |
46-
| 2025-10-06 08:00 | 2025-10-06 18:15 | 1.2W | MD16203 | LRBB Beta-beat compensation with flat-optics | [Shift Start](logbook://2025-10-06,1081,4401590) / [Summary](logbook://2025-10-06,1081,4402045) ([OP][op_MD16203]{.cern_login target=_blank}) |
47+
| 2025-10-06 08:00 | 2025-10-06 18:15 | 1.2W | MD16203 | LRBB Beta-beat compensation with flat-optics | [Shift Start](logbook://2025-10-06,1081,4401590) / [Summary](logbook://2025-10-06,1081,4402045) ([OP][op_MD16203]{.cern_login target=_blank}) |
48+
| 2025-10-28 12:30 | 2025-10-28 22:00 | 0.5W 0.5WN | Commissioning | Pb Ion Optics Part 4 | [Shift Start](logbook://2025-10-28,1081,4418083) / [Summary](logbook://2025-10-28,1081,4421639) ([OP][op_ionpart3]{.cern_login target=_blank}) |
49+
| 2025-11-02 23:00 | 2025-11-02 23:59 | 0.15HN | END-of-FILL | Luminosity measurement while trimming waist | [Shift Start](logbook://2025-11-02,1081,4435672) / [Summary](logbook://2025-11-02,1081,4441388) ([OP][op_lumitest]{.cern_login target=_blank}) |
50+
| 2025-11-04 22:30 | 2025-11-05 03:40 | 0.5WN | MD17443 | ADT-ACD measurements in OP config (indiv,MO,Q') | [Shift Start](logbook://2025-11-04,1081,4439816) / [Summary](logbook://2025-11-05,1081,4440023) ([OP][op_opconfig]{.cern_login target=_blank}) |
51+
| 2025-11-06 03:00 | 2025-11-06 09:30 | 0.8WN | MD17383 | New IR7 optics, squeeze validation | [Shift Start](logbook://2025-11-06,1081,4441445) / [Summary](logbook://2025-11-06,1081,4441726) ([OP][op_ir7squeeze]{.cern_login target=_blank}) |
52+
| 2025-11-07 00:01 | 2025-11-07 11:30 | 1WN 0.25W | MD16906 | Extra squeeze steps | [Shift Start](logbook://2025-11-07,1081,4442872) / [Summary](logbook://2025-11-07,1081,4443242) ([OP][op_md16906]{.cern_login target=_blank}) |
53+
| 2025-11-07 11:30 | 2025-11-07 13:30 | 0.25W | MD17423 | Identifying beam energy discrepancy | [Shift Start](logbook://2025-11-07,1081,4442872) / [Summary](logbook://2025-11-07,1081,4443242) |
54+
| 2025-11-08 00:07 | 2025-11-08 11:00 | 0.5H | Commissioning | PbPb Ion Optics Part 5 | [Shift Start](logbook://2025-11-08,1081,4444644) / [Summary](logbook://2025-11-08,1081,4444815) ([OP][op_pbpart3]{.cern_login target=_blank}) |
4755
<!-- Logbook Links: [LINK_NAME](logbook://date, logbook_id, event_id) -->
4856

4957
<!-- OP logbook links below -->
@@ -58,6 +66,7 @@ The tables below can be sorted by clicking next to the column headers.
5866
[op_global_corrs_60cm_18cm]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4243717
5967
[op_validation_everything]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4245973
6068
[op_ion_optics_part1]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4258231
69+
[op_vdm]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4292225
6170
[op_ion_optics_part2]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4295553
6271
[op_po]:https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4315094
6372
[op_hierarchy_breaking_injection]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4315806
@@ -68,7 +77,12 @@ The tables below can be sorted by clicking next to the column headers.
6877
[op_MD15003]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4378082
6978
[op_MD15464]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4380521
7079
[op_MD16203]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4402046
71-
80+
[op_ionpart3]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4420108
81+
[op_lumitest]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4436213
82+
[op_opconfig]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4440024
83+
[op_ir7squeeze]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4441728
84+
[op_md16906]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4443198
85+
[op_pbpart3]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4444835
7286
<!-- Tooltips -->
7387

7488
*[Injection Optics Global Corrs]: Test of tools. Rephased BPMs. Confirmed local coupling corrections, measured virgin machine at injection optics and did global corrections. Reached similar performance as previous years. First attempts of total phase corrections, used the ones from 2024. First use of new Segment-by-Segment.
@@ -94,5 +108,8 @@ The tables below can be sorted by clicking next to the column headers.
94108
*[Refining optics corrections at top energy]: Local chroma at injection wihtout QFB (last point with QFB). ADT-ACD settup in optics mode at injection, 10mm kicks. ADT-ACD at FT, large enough kick to measure linear optics. At FT had problem with BPM concentrator for B1. Local chroma of B2 at 18cm. Kmod at 160 and 120 urad for effect of crossing-scheme on lumi imbalance. Kmod at 120urad with Felix waist knob. Kmod and ACD with 28april waist knob. ACDkicks and xing scan for b6 feed-down to f0140 RDT.
95109
*[Pb Ion Optics Part 3]: Attempted to validate Pb optics with enf of squeeze corrections from last time, as filler during validation of loss maps during vaccuum issue. Quick measurement at injection to check BPMs, but then OP were unable to get cycle to ramp due to inconsistency in ADT settings. Gave up after several hours as LHC switching back to lumi production. Includes summary of previous ion fill results.
96110
*[LRBB Beta-beat compensation with flat-optics]: Tried for the first time to take measurements of the LRBB-optics in both beam-1 and beam-2, using 2 separate fills. No issues encountered when switching the role of weak/strong beams between beam1/2. In Beam1 there was no visible effect on the lifetime. In beam2 however the lifetime appeared to degrade slightly as the MO were removed.
111+
*[New IR7 optics, squeeze validation]: Checked the new IR7 optics along the squeeze down to 60cm/18cm, including new corrections scaled from nominal cycle. Beta-beating and normalised dispersion good at all points in the squeeze. Vertical dispersion bump observed in IR7, getting stronger as the squeeze progresses (up to Dy=1m). No issue for the second fill of MD with lossmaps but should be corrected for operational use. The sdisp knob is to be included in model. Kmodulation at 18cm shows similar results to nominal ion cycle.
112+
*[Extra squeeze steps]: First part, tested a new squeeze down to 40cm/12cm. Stopped at 50cm/15cm where the optics were very good (similar to 60cm/18cm). At 40cm/12cm a peak beat-beating of 20% was observed (acceptable, should be corrected) and small waist shift. Unforeseen negative effects of Stephane's correction knobs brought peak beta-beating to 100% in the arcs and betastar to ~7-9cm. Did not proceed with fill of nominal bunches.
113+
*[Identifying beam energy discrepancy]: Benchmark data at 40cm/12cm with various DP_TRIM_PERMIL values to be analyzed offline.
97114

98115
--8<-- "docs/logbook/footer_shifts"

docs/logbook/SPS/2025_sps.md

Lines changed: 1 addition & 0 deletions
Original file line numberDiff line numberDiff line change
@@ -18,6 +18,7 @@
1818
| 2025-09-17 15:00 | 2025-09-17 18:00 | 0.5W | MD | Panos/Ewen | Tests of CC measurement procedures | [Start](logbook://2025-09-17,2621,4389546) |
1919
| 2025-09-19 09:00 | 2025-09-19 13:00 | 0.5W | MD | Panos | Tests of CC measurement procedures | [Start](logbook://2025-09-19,2621,4390955) |
2020
| 2025-09-23 09:00 | 2025-09-23 13:00 | 0.5W | MD | Panos/Mattia | Tests of CC measurement procedures | [Start](logbook://2025-09-23,2621,4393166) |
21+
| 2025-10-17 10:00 | 2025-10-17 14:00 | 0.5W | MD | Panos/Ewen | Tests of CC measurement procedures | [Start](logbook://2025-10-17,2621,4411175) |
2122

2223
<!-- Tooltips -->
2324

0 commit comments

Comments
 (0)